OpenStudy (anonymous):

Given 2 power x = 3 power y = 6 power p , show that p = xy/ x+y. Obviously this is a logarithm question.

OpenStudy (aravindg):

so the question is ' \[\large 2^x=3^y=6^p\]

OpenStudy (anonymous):

noted

OpenStudy (aravindg):

ok taking logarithm \[x \log 2=y \log 3=p \log 6\]

OpenStudy (anonymous):

yea right that was what you told me to learn which is the basics i did that

OpenStudy (sirm3d):

logarithm as the required solution?

OpenStudy (anonymous):

no show that p = xy/ x+y

OpenStudy (aravindg):

now log(6)=log 3+log 2

OpenStudy (anonymous):

understood that.

OpenStudy (anonymous):

The confusing part should be somewhere after this steps.

OpenStudy (anonymous):

which i'm stuck.

OpenStudy (aravindg):

wait a sec pls i am doing on paper

OpenStudy (aravindg):

@Sgstudent i also got stuck :( ...lets see if sirm3d can help :)

OpenStudy (sirm3d):

i'll post a non-logarithm solution. \[\Large{2^x=6^p\\(2^x)^y=(6^p)^y\\2^{xy}=y^{py}}\]

OpenStudy (sirm3d):

that's \[\Large 2^{xy}=6^{py}\]

OpenStudy (anonymous):

lol i need to use logarithm to solve this type of question is going to be out in my upcoming class test this thursday.

OpenStudy (aravindg):

ok got it !!

OpenStudy (aravindg):

:)

OpenStudy (sirm3d):

ahhh. you wan't logarithm solution. \[x\log2=p\log6\\xy\log2 = py\log 6\\y\log 3 = p\log 6\\xy\log 3 = px\log 6\]

OpenStudy (sirm3d):

add the equations: \[\Large xy\log 2 + xy\log 3 = py\log 6 + px\log 6\]

OpenStudy (aravindg):

from my 2nd step \[\large \dfrac{y}{x}=\dfrac{\log 2}{\log 3}\]

OpenStudy (anonymous):

I will review both solutions and see which i can understand easier.

OpenStudy (sirm3d):

\[\Large {xy\log 6 = \log 6 (py + px)\\xy=p(y+x)\\\frac{xy}{x+y}=p}\]

OpenStudy (anonymous):

oh sirm3d can you compile all your workings into one?

OpenStudy (anonymous):

easier to refer to thanks

OpenStudy (sirm3d):

ok

OpenStudy (anonymous):

i think AravindG method will be much longer.

OpenStudy (anonymous):

Thank you for the effort.

OpenStudy (aravindg):

now \[p(\log3+\log 2)=y \log 3 \] divide by \(\log 3 +\log 2 \) on both sides \[p=\dfrac{y \log3}{\log 3+\log 2}\] divide by log 3 on both numerator and denominator \[p=\dfrac{y}{{1+ (\dfrac{\log 2}{\log 3}}}\]

OpenStudy (aravindg):

substitute \[\dfrac{\log 2 }{\log 3}=\dfrac{y}{x}\] and yu will get the answer :)

OpenStudy (anonymous):

lol i'm amazed

OpenStudy (aravindg):

amazed?

OpenStudy (anonymous):

this question took my class about 50 mins to solve and came out with the solution you mentioned/

OpenStudy (sirm3d):

1st equation\[\Large{2^x=6^p\\x\log2 = p\log6}\]multiply both sides by y\[\Large xy\log 2 = py\log 6\] 2nd equation:\[\Large {3^y=6^p\\y\log3=p\log6\\xy\log 3 = px\log 6}\] add the two equations\[\Large{xy\log 2 + xy\log 3 = py\log 6 + py\log 6\\xy(\log 2 + \log 3)=p(x+y)\log 6\\xy\log 6=p(x+y)\log 6 \\\frac{xy\log 6}{(x+y)\log 6}=p\\\frac{xy}{x+y}=p}\]

OpenStudy (aravindg):

@Sgstudent who mentioned ?

OpenStudy (anonymous):

you

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

but i did not copy down the solution i prefer something simpler

OpenStudy (anonymous):

hence, i decided to come here and look for the solution i want :P

OpenStudy (sirm3d):

now you have two different solutions. ;-)

OpenStudy (aravindg):

:) all the best ..happy to know that you learned the basics :)

OpenStudy (anonymous):

thanks both of you

OpenStudy (aravindg):

yw :)

OpenStudy (anonymous):

i know why you multiply y in the first equation it is because p(x+y) = px+py am i right?

OpenStudy (anonymous):

this is called deduction method i suppose

OpenStudy (sirm3d):

i need to add xlog 2 and y log 3 and combine as one term. hence it can only be done if they have the same literal coefficients, xy.

OpenStudy (anonymous):

noted.

OpenStudy (anonymous):

AravindG

OpenStudy (anonymous):

please help if possible :)

OpenStudy (anonymous):

I need some tips on delivering a speech without a script etc.